OpenStudy (anonymous):

(CAN SOMEONE CHECK MY ANSWER PLS, WILL GIVE MEDAL) A tuition scale for a state university shows an increase of $1,100 each year. If the initial cost was $27,000. What will the cost of tuition be in twenty two years?

OpenStudy (anonymous):

Each year you are adding $1,100 Tuition is $27,000, $28,100, $29,200, $30,300 ... It forms an arithmetic sequence An = A1 +d(n-1) A1 = 27,000 d = 1100 n=22 Tuition in 22 years = 27,000 + 1100(22-1) Tuition in 22 years=$50,100

OpenStudy (anonymous):

Can someone just tell me if I'm right or wrong?

OpenStudy (paperbacon):

Correct. It is an arithmetic sequence. That is the right answer

OpenStudy (anonymous):

Thank you so much!